# SlideChat: A Large Vision-Language Assistant for Whole-Slide Pathology Image Understanding

### v2

@@ Abstract
Despite the progress made by multimodal large language models (MLLMs) in computational pathology, they remain limited by a predominant focus on patch-level analysis, missing essential contextual information at the whole-slide level. The lack of large-scale instruction datasets and the gigapixel scale of whole slide images (WSIs) pose significant developmental challenges. In this paper, we present SlideChat, the first vision-language assistant capable of understanding gigapixel whole-slide images, exhibiting excellent multimodal conversational capability and response complex instruction across diverse pathology scenarios. To support its development, we created SlideInstruction, the largest instruction-following dataset for WSIs consisting of 4.2K WSI captions and 176K VQA pairs with multiple categories. Furthermore, we propose SlideBench, a multimodal benchmark that incorporates captioning and VQA tasks to assess SlideChat’s capabilities in varied clinical settings such as microscopy, diagnosis. Compared to both general and specialized MLLMs, SlideChat exhibits exceptional capabilities, achieving state-of-the-art performance on 18 of 22 tasks. For example, it achieved an overall accuracy of 81.17% on SlideBench-VQA (TCGA), and 54.15% on SlideBench-VQA (BCNB). We will fully release SlideChat, SlideInstruction and SlideBench as open-source resources to facilitate research and development in computational pathology.
@@ 1 Introduction
Computational pathology aims to improve the analysis of digitized tissue samples, such as whole slide images (WSIs), by applying artificial intelligence to aid in the diagnosis, identification, and understanding of disease (Song et al., 2023). Recently, the development of this field has gained rapid momentum, mainly driven by breakthroughs in the visual foundation model (Chen et al., 2024b; Xu et al., 2024a; Vorontsov et al., 2024). These models learn generalized representations by pre-training on large-scale data and perform well in various downstream tasks, including rare cancer detection and biomarker prediction. Building on this base, integration with the powerful Large Language Models (LLMs) further advances the development of the Multimodal Large Language Model (MLLMs) (Lu et al., 2024b), which has made great strides in responding to more complex, open-ended visual queries, enabling it to serve as a versatile assistant at various stages of medical care, including clinical decision-making, education, and research (see Figure 1).
Nevertheless, there are three major challenges that hinder the development and use of pathology MLLMs for real-world clinical applications. First, it is challenging to develop a MLLMs architecture that can effectively capable of gigapixel whole slides (e.g., 100,000 × 100,000 pixels). Existing models  (Lu et al., 2024b; Sun et al., 2024; Seyfioglu et al., 2024) often process whole slides by extracting small patch/ROI-level data for subsequent analysis, resulting in limited understanding of global slide context and suboptimal performance in some complex pathological analysis. Second, publicly available multi-modal pathology slide dataset are relatively scarce and of varying quality (Guo et al., 2024; Chen et al., 2023; 2024a), which limits the development of MLLMs trained on such data. Third, current pathology MLLMs (Lu et al., 2024b) are developed using closed-source data from medical institutions. Consequently, the model weights and associated instructional datasets are typically not made full publicly available, thereby restricting their broader applicability in clinical research and applications.
@@
Figure 1: SlideChat’ is the first large vision-language assistant specifically designed for whole-slide pathology analysis. SlideChat can generates comprehensive descriptions of whole-slide images and provides contextually relevant responses across various applications.
@@
In this paper, we present SlideChat, the first open-source vision-language assistant capable of understanding gigapixel whole-slide images. First, SlideChat is trained on SlideInstruction, a large-scale multi-modal instruction dataset encompassing data from The Cancer Genome Atlas (TCGA) (Hutter & Zenklusen, 2018) via our specifically designed data processing pipeline (see Figure 3 (A)). SlideInstruction contains 4,181 WSI-caption pairs and 175,754 visual question-answer pairs from 3,294 patients, covering 10 cancer types. The question-answer pairs include both open-ended and closed-ended questions, further divided into 13 subcategories, covering a diverse range of clinical tasks such as tumor grading. SlideInstruction is more than 20 times larger than previous public instruction datasets in the number of instructions(see Table 11 in Appendix). Second, we propose SlideChat, a novel architecture in LLaVA style for capable multi-modal analysis of gigapixel whole slides. As shown in Figure 2, a gigapixel whole slide is first divided into a series of patches, each of which is individually processed by a patch-level encoder to extract local features. The resulting long sequence of feature tokens is then processed by a slide-level encoder employing sparse attention to aggregate the slide-level features. Finally, the aggregated features are fed into a Large Language Model via a projector, which processes user queries and generates responses.
@@
To systematically evaluate the performance of SlideChat in real-world scenarios, we establish a comprehensive digital pathology benchmark (see Table 1) encompassing more than 20 clinical tasks, using data from both TCGA and the in-the-wild Early Breast Cancer Core-Needle Biopsy (BCNB) dataset. This resulted in three test sets: SlideBench-Caption, comprising 734 WSI-caption pairs; SlideBench-VQA (TCGA), comprising 7,827 VQA pairs covering 13 different tasks; and SlideBench-VQA (BCNB), including a total of 7,274 VQA entries from 1,058 patients, covering seven different tasks. Additionally, we compare the performance of SlideChat on another externally proposed dataset, WSI-VQA (Chen et al., 2024a), to further validate its effectiveness. We compare our model with the currently available state-of-the-art general and medical-specialized MLLMs including GPT-4o, LLava-Med (Li et al., 2024a), MedDr (He et al., 2024). Benefiting from large-scale, high-quality training and effective local-global context modelling, SlideChat achieves state-of-the-art performance on 18 out of 22 tasks, with significant improvements over the second-best method 10% on 9 tasks on four benchmarks. Specifically, SlideChat achieves an average accuracy improvement of 13.47% over the second-best model on SlideBench-VQA (TCGA), an average improvement of 12.71% on SlideBench-VQA (BCNB), and an improvement of 5.82% on WSI-VQA. Finally, to accelerate research progress in digital pathology, we make SlideChat fully open-weight, including source code and model weights as well as instruction and benchmark dataset. The key contributions are summarized four-fold in the following:
@@
  • We establish SlideBench, a WSIs multi-modal benchmark comprising SlideBench-Caption, SlideBench-VQA (TCGA), and SlideBench-VQA (BCNB), covering 21 different clinical tasks.
  • We will release SlideChat, SlideInstruction and SlideBench as open-source resources to facilitate research and development in computational pathology.
@@ 2 Related Works
Whole slide images are pivotal in modern pathology, enabling comprehensive analysis of tissue samples for tasks such as predicting patient prognosis, classifying cancer subtypes, and identifying biomarkers (Song et al., 2023; Shao et al., 2023; Li et al., 2024b; Spronck et al., 2023). Recent studies have leveraged pathology foundational models (Wang et al., 2024; Ahmed et al., 2024; Xu et al., 2024b) to enhance WSIs analysis, either through fine-tuning for specific downstream tasks or by employing zero-shot prediction approaches in CLIP (Radford et al., 2021) style. Although these models are effective in task-specific applications, their reliance on fine-tuning or limited zero-shot capabilities restricts their generalizability across diverse and complex user instructions.
@@
MLLMs in Computational Pathology
The paradigm of MLLMs enables to effectively respond to more complex, open-ended visual queries while processing pathology image, thus providing significant value across various medical stages. PathChat (Lu et al., 2024b) is a vision-language assistant designed for pathology, developed with 450K private instruction pairs to handle both visual and natural language queries. QuiltInstruct (Seyfioglu et al., 2024) is a large-scale dataset comprising 107K question-answer pairs. Building on QuiltInstruct, Quilt-LLAVA (Seyfioglu et al., 2024) is a model designed for diagnostic reasoning across multiple image patches, leveraging its extensive question-answer pairs to accurately interpret complex H&E data. PathAsst (Sun et al., 2024) combines a pathology-specific CLIP model with Vicuna-13b (Chiang et al., 2023) to create a multimodal generative foundational model tailored for pathology. However, current MLLMs primarily focus on patch or region-of-interest (ROI) data, limiting their utility for slide-level clinical applications where broader contextual understanding is crucial.
@@ 3 SlideChat
Figure 3: (A) Overview of the SlideInstruction generation pipeline. We prompt GPT-4 to extract the WSI-Caption, Open-set VQA and Closed-set VQA from reports. (B) For the generated Closed-set VQA, we employ LLMs to filter low-quality QA pairs and involve pathologists for validation, resulting in the creation of SlideBench-VQA. (C) Examples of WSI caption and instruction-following scenarios in microscopy, diagnostics, and clinical applications. For additional examples, please refer to  Figure 6 in the Appendix.
@@ 3.1 Architecture
To achieve the goal of analyzing gigapixel whole-slide images in a multimodal setting, as shown in Figure 2, SlideChat consists of four key designs: the patch-level encoder, the slide-level encoder, the multimodal projector module, and the large language model. Our method starts by partitioning the WSI into smaller 224 × 224 pixel patches, making it computationally feasible to process such large images. These patches are then passed through a well-trained, frozen patch-level encoder (Lu et al., 2024a), which extracts localized features from each individual patch, capturing fine-grained details such as cellular structures. Building on this, we employ LongNet (Ding et al., 2023; Xu et al., 2024a) as slide-level encoder to enhance the patch-level embeddings and capture global patterns across the entire slide. This encoder uses sparse attention mechanisms to aggregate both local and global contextual information, enabling the model to perceive intricate local features while capturing the broader context, which is critical for comprehensive pathological assessments. Following the slide-level encoding, SlideChat incorporates a multimodal projection layer that maps these aggregated visual features into a unified space aligned with the LLM. This ensures that the visual features extracted from the WSIs are effectively transformed into representations compatible with the language model, facilitating seamless integration and interaction between visual and textual data. Concurrently, the model accepts natural language instructions from users, such as “What is the type of tumor in the image?”. These textual queries are processed by the LLM, which comprehends the textual input and integrates it with the visual features extracted from the WSIs, enabling accurate and contextually relevant diagnostic responses. This multimodal reasoning capability allows SlideChat to provide accurate and contextually relevant answers to complex pathology-related questions, thereby supporting clinical decision-making, education, and research across various medical stages.
@@ 3.2 Data
There is a notable lack of large-scale multimodal pathology datasets supporting the training of vision-language assistants for whole-slide image understanding. To support the training of SlideChat, we develop SlideInstruction, a comprehensive instruction dataset, sourced from the TCGA database, comprising 4,915 whole slide image (WSI)-report pairs from 4,028 patients. Figure 3 illustrates our entire data curation pipeine. We initially prompt GPT-4 to refine the pathology reports, clean up the noise in the report including unrelated symbols, technical details of pathology department procedures, specimen handling and processing information, redundant administrative or legal statements, and some repeated information. For the refined pathology reports, we further employ GPT-4 to generate high-quality multimodal data, comprising two main components: (1) WSI-Caption Data: We craft concise, clinically relevant summaries for each whole slide image by prompting the language model to extract key pathological findings. These summaries were structured into coherent paragraphs that highlighted crucial clinical details such as diagnostic results, tumor characteristics, margin status, and lymph node involvement, ensuring the caption dataset is both focused and informative. (2) WSI Instruction-Following Data: To enhance the model’s ability to follow instructions and improve its comprehension of pathology images, we leveraged GPT-4 to generate tailored question-and-answer pairs for each WSI report. Drawing inspiration by PathChat (Lu et al., 2024b), we structure these questions into three “broad” categories—microscopy, diagnosis, and clinical considerations—which represent key stages in the pathology workflow, and thirteen “narrow” categories focusing on specific aspects within each stage (Figure 1 B). Our carefully crafted prompts are detailed in Section A.2.2. To create a comprehensive instructional dataset, we generated two open-ended and two closed-ended QA pairs within each narrow category for every WSI report. Regarding the train/test split, it is worth noting that the WSI-report datasets from TCGA includes two types: (a) one report linked to multiple WSIs, and (b) one report linked to a single WSI. For type (a), where specific diagnostic details may not align perfectly with each WSI, we include all WSIs in the training set to introduce some “noisy data”, which can enhance model robustness. For type (b), 80% of WSIs are allocated to the training set and 20% to the test set. Finally, there are 4,181 WSIs for training and 734 WSIs for testing. Consequently, we construct a large-scale training set named SlideInstruction, comprising 4,181 WSI captions and 175,753 instruction-following VQA pairs across various broad and narrow categories.
Table 1: Statistical information of SlideBench.
Subset #Patient #Data #Tumor #Tasks Answer Type Evaluation Metirc
SlideBench-Caption 734 734 10 1 Free From BLEU, Rouge, GPT score
SlideBench-VQA (TCGA) 732 7,827 10 13 A/B/C/D Accuracy
SlideBench-VQA (BCNB) 1058 7,274 1 7 A/B/C/D Accuracy
@@
To systematically evaluate the performance of SlideChat, We incorporate the remaining 734 WSI captions along with a substantial number of closed-set VQA pairs to establish evaluation benchmark. First, we construct a test set named SlideBench-Caption based on the WSI-Caption data to evaluate the model’s ability to generate accurate and coherent descriptions of whole slide images. Secondly, we construct SlideBench-VQA (TCGA) based on closed-set visual question-answering (VQA) pairs along with test WSIs, aiming to evaluate various aspects of model performance. As shown in Figure 3 (B), to improve the quality of the testing benchmarks, we employ four advanced large language models, including GPT-4 (Achiam et al., 2023), InternLM2-Chat-7B (Cai et al., 2024), Qwen-7B-Chat (Bai et al., 2023), and DeepSeek-7B-Chat, to filter closed-set VQAs by predicting answers based solely on the question text. Any questions for which at least three of these models provided correct answers are subsequently excluded. Following this automated filtering, five expert pathologists are invited to review and amend the remaining questions. The review process are guided by the following criteria: (1) Whether the correct answer necessitates image interpretation; (2) Whether the question and its corresponding answer are logically and coherently structured; and (3) Whether the question aligns appropriately with the designated broad and narrow categories. QA pairs failing to meet these criteria are excluded by the pathologists. Consequently, the SlideBench-VQA (TCGA) comprises 7,827 VQAs across 13 categories, with some examples illustrated in Figure 3 C. Additionally, we incorporate the in-the-wild Early Breast Cancer Core-Needle Biopsy (BCNB) WSI dataset (Xu et al., 2021), which encompasses a diverse patient population and a variety of clinical task labels, to enhance the test set benchmark and more comprehensively assess the model’s generalization capabilities. In detail, we convert the BCNB dataset into a VQA format by rephrasing the classification objectives into a specific template as questions, while transforming the original multi-class labels into selectable options, and integrate it into SlideBench as an external subset, named SlideBench-VQA (BCNB). This dataset comprises 7,247 VQA pairs from 1,058 patients, specifically designed to evaluate SlideChat’s zero-shot generalization capability across 7 distinct classification tasks. More detailed information about SlideBench is provided in Table 1.

### v3

@@ Abstract
Despite the progress made by multimodal large language models (MLLMs) in computational pathology, they remain limited by a predominant focus on patch-level analysis, missing essential contextual information at the whole-slide level. The lack of large-scale instruction datasets and the gigapixel scale of whole slide images (WSIs) pose significant developmental challenges. In this paper, we present SlideChat, the first vision-language assistant capable of understanding gigapixel whole-slide images, exhibiting excellent multimodal conversational capability and response complex instruction across diverse pathology scenarios. To support its development, we created SlideInstruction, the largest instruction-following dataset for WSIs consisting of 4.2K WSI captions and 176K VQA pairs with multiple categories. Furthermore, we propose SlideBench, a multimodal benchmark that incorporates captioning and VQA tasks to assess SlideChat’s capabilities in various settings such as microscopy, diagnosis and clinical. Compared to both general and specialized MLLMs, SlideChat exhibits exceptional capabilities, achieving state-of-the-art performance on 18 of 22 tasks. For example, it achieved an overall accuracy of 81.17% on SlideBench-VQA (TCGA), and 54.15% on SlideBench-VQA (BCNB). Our code, data, and model is publicly accessible at https://uni-medical.github.io/SlideChat.github.io.
@@ 1 Introduction
Computational pathology aims to improve the analysis of digitized tissue samples [15, 25], such as whole slide images (WSIs), by applying artificial intelligence to aid in the diagnosis, identification, and understanding of disease. Recently, the development of this field has gained rapid momentum, mainly driven by breakthroughs in the visual foundation model [9, 31, 28]. These models learn generalized representations by pre-training on large-scale data and perform well in various downstream tasks, such as cancer detection and biomarker prediction. Building on this base, integration with the powerful Large Language Models (LLMs) further advances the development of the Multimodal Large Language Model (MLLMs) [21], which has made great strides in responding to more complex, open-ended visual queries, enabling it to serve as a versatile assistant at various stages of medical care, including clinical decision-making, education, and research.
@@
Figure 1: SlideChat is the first large vision-language assistant specifically designed for whole-slide pathology analysis. SlideChat can generate comprehensive descriptions of whole-slide images and provide contextually relevant responses across various applications.
Nevertheless, there are some major challenges that hinder the development and use of pathology MLLMs for real-world applications. First, it is challenging to develop a MLLM architecture that can effectively be capable of gigapixel whole slides (e.g., 100,000 × 100,000 pixels). Existing models  [21, 27, 23] often process whole slides by extracting small patch/ROI-level data for subsequent analysis, resulting in a limited understanding of global slide context and suboptimal performance in some complex pathological analysis. Second, publicly available multi-modal pathology slide datasets are relatively scarce and of varying quality [13, 7, 8], which limits the development of MLLMs trained on such data.
@@
In this paper, we present SlideChat (see Fig. 1), the first open-source vision-language assistant capable of understanding gigapixel whole-slide images. First, SlideChat is trained on SlideInstruction, a large-scale multi-modal instruction dataset encompassing data from The Cancer Genome Atlas (TCGA) [16] via our specifically designed data processing pipeline (see Fig. 3 (A)). SlideInstruction contains 4,181 WSI-caption pairs and 175,753 visual question-answer pairs from 3,294 patients, covering 10 cancer types. The question-answer pairs include both open-ended and closed-ended questions, further divided into 13 subcategories, covering a diverse range of clinical tasks. SlideInstruction is more than 20 times larger than previous public instruction datasets in the number of instructions. Second, we propose SlideChat, a novel architecture in LLaVA style for capable multi-modal analysis of gigapixel whole slides. A gigapixel whole slide is first divided into a series of patches, each of which is individually processed by a patch-level encoder to extract local features. The resulting long sequence of patch tokens is then processed by a slide-level encoder employing sparse attention to enhance the slide-level features. Finally, enhanced patch tokens are fed into the LLM via a projector, which processes user queries and generates responses.
@@
To systematically evaluate the performance of SlideChat in real-world scenarios, we establish a comprehensive digital pathology benchmark named SlideBench, encompassing more than 20 clinical tasks, using data from both TCGA and the in-the-wild Early Breast Cancer Core-Needle Biopsy (BCNB) dataset. This resulted in three test sets: SlideBench-Caption, comprising 734 WSI-caption pairs; SlideBench-VQA (TCGA), comprising 7,827 VQA pairs covering 13 different tasks; and SlideBench-VQA (BCNB), including a total of 7,274 VQA entries from 1,058 patients, covering seven different tasks. Additionally, we compare the performance of SlideChat on another externally proposed dataset, WSI-VQA [8], to further validate its effectiveness. We compare our model with the currently available state-of-the-art general and medical-specialized MLLMs including GPT-4o, LLava-Med [17], MedDr [14] and Quilt-LLaVA [23]. Benefiting from large-scale, high-quality training and effective local-global context modeling, SlideChat achieves state-of-the-art performance on 18 out of 22 tasks, with significant improvements over the second-best method 10% on 9 tasks on four benchmarks. Specifically, SlideChat achieves an average accuracy improvement of 13.47% over the second-best model on SlideBench-VQA (TCGA), an average improvement of 12.59% on SlideBench-VQA (BCNB), and an improvement of 5.82% on WSI-VQA. Finally, to accelerate research progress in digital pathology, we make SlideChat fully open-weight, including source code and model weights as well as instruction and benchmark datasets. The key contributions are summarized four-fold in the following:
@@
  • We establish SlideBench, a WSI multi-modal benchmark comprising SlideBench-Caption, SlideBench-VQA (TCGA), and SlideBench-VQA (BCNB), covering 21 different clinical tasks.
  • We release SlideChat, SlideInstruction and SlideBench as open-source resources to facilitate research and development in computational pathology.
@@ 2 Related Works
Whole slide images are pivotal in modern pathology, enabling comprehensive analysis of tissue samples for tasks such as predicting patient prognosis, classifying cancer subtypes, and identifying biomarkers [25, 24, 18, 26, 19, 35]. Recent studies have leveraged pathology foundational models [29, 3, 32] to enhance WSIs analysis, either through fine-tuning for specific downstream tasks or by employing zero-shot prediction approaches in CLIP [22] style. Although these models are effective in task-specific applications, their reliance on fine-tuning or limited zero-shot capabilities restricts their generalizability across diverse and complex user instructions.
MLLMs in Computational Pathology
The paradigm of MLLMs enables them to effectively respond to more complex, open-ended visual queries while processing pathology images, thus providing significant value across various medical stages. PathChat [21] is a vision-language assistant designed for pathology, developed with 450K private instruction pairs to handle both visual and natural language queries. Quilt-Instruct [23] is a large-scale dataset comprising 107K question-answer pairs. Building on Quilt-Instruct, Quilt-LLAVA [23] is a model designed for diagnostic reasoning across multiple image patches, leveraging its extensive question-answer pairs to accurately interpret complex H&E images. PathAsst [27] combines a pathology-specific CLIP model with Vicuna-13b [10] to create a multimodal generative foundational model tailored for pathology. However, current MLLMs primarily only focus on patch or ROI data, limiting their utility for slide-level clinical applications where broader contextual understanding is crucial.
@@ 3.1 Architecture
To achieve the goal of analyzing gigapixel whole-slide images in a multimodal setting, as shown in Fig. 2, SlideChat consists of four key designs: the patch-level encoder, the slide-level encoder, the multimodal projector module, and the large language model. Our method starts by partitioning the WSI into smaller 224 × 224 pixel patches, making it computationally feasible to process such large images. These patches are then passed through a well-trained, frozen patch-level encoder [20], which extracts localized features from each individual patch, capturing fine-grained details such as cellular structures. Building on this, we employ LongNet [12, 31] as slide-level encoder to enhance the patch-level embeddings and capture global patterns across the entire slide. This encoder uses sparse attention mechanisms to learn both local and global contextual information, enabling the model to perceive intricate local features while capturing the broader context, which is critical for comprehensive pathological assessments. Following the slide-level encoding, SlideChat incorporates a multimodal projector that maps these patch tokens into a unified space aligned with the LLM. This ensures that the visual features extracted from the WSIs are effectively transformed into representations compatible with the language model, facilitating seamless integration and interaction between visual and textual data. Concurrently, the model accepts natural language instructions from users, such as “What is the type of tumor in the image?”. These textual queries are processed by the LLM, which comprehends the textual input and integrates it with the visual features extracted from the WSIs, enabling accurate and contextually relevant diagnostic responses. This multimodal reasoning capability allows SlideChat to provide accurate and contextually relevant answers to complex pathology-related questions, thereby supporting clinical decision-making, education, and research across various medical stages.
Figure 3: (A) Overview of the SlideInstruction generation pipeline. We prompt GPT-4 to extract the WSI-Caption, Open-set VQA and Closed-set VQA from reports. (B) For the generated Closed-set VQA, we employ LLMs to filter low-quality QA pairs and involve pathologists for validation, resulting in the creation of SlideBench-VQA. (C) Examples of WSI caption and instruction-following scenarios in microscopy, diagnostics, and clinical applications. For additional examples, please refer to Supplementary Material.
@@ 3.2 Data
There is a notable lack of large-scale multimodal pathology datasets supporting the training of vision-language assistants for whole-slide image understanding. To support the training of SlideChat, we develop SlideInstruction, a comprehensive instruction dataset, sourced from the TCGA database, comprising 4,915 WSI-report pairs from 4,028 patients. Fig. 3 illustrates our entire data curation pipeine. We initially prompt GPT-4 to refine the pathology reports, and clean up the noise in the report including unrelated symbols, technical details of pathology department procedures, specimen handling and processing information, redundant administrative or legal statements, and some repeated information. For the refined pathology reports, we further employ GPT-4 to generate high-quality multimodal data, comprising two main components: (1) WSI-Caption Data: We craft concise, clinically relevant summaries for each whole slide image by prompting the language model to extract key pathological findings. These summaries were structured into coherent paragraphs that highlighted crucial clinical details such as diagnostic results, tumor characteristics, margin status, and lymph node involvement, ensuring the caption dataset is both focused and informative. (2) WSI Instruction-Following Data: To enhance the model’s ability to follow instructions and improve its comprehension of pathology images, we leveraged GPT-4 to generate tailored question-and-answer pairs for each WSI report. Drawing inspiration from PathChat [21], we structure these questions into 3 broad categories—microscopy, diagnosis, and clinical considerations—which represent key stages in the pathology workflow, and 13 narrow categories focusing on specific aspects within each stage (Fig. 1 B). To construct a comprehensive instructional dataset, we began by generating two open-ended and two closed-ended QA pairs for each report within each specific category, taking advantage of the rich content found in the majority of TCGA reports. Then, some QA pairs were excluded, such as those lacking corresponding category information. Regarding the train/test split, it is worth noting that the WSI-report datasets from TCGA include 2 types: (a) one report linked to multiple WSIs, and (b) one report linked to a single WSI. For type (a), where specific diagnostic details may not align perfectly with each WSI, we include all WSIs in the training set to introduce some “noisy data”, which can enhance model robustness. For type (b), 80% of WSIs are allocated to the training set and 20% to the test set. Finally, there are 4,181 WSIs for training and 734 WSIs for testing. Consequently, we construct a large-scale training set named SlideInstruction, comprising 4,181 WSI captions and 175,753 instruction-following VQA pairs across 13 narrow categories.
@@
To systematically evaluate the performance of SlideChat, we include the remaining 734 WSI captions from the test set, along with a substantial number of closed-set VQA pairs, to establish a comprehensive evaluation benchmark. First, we construct a set named SlideBench-Caption based on the WSI-Caption data to evaluate the model’s ability to generate accurate and coherent descriptions of whole slide images. Secondly, we construct SlideBench-VQA (TCGA) based on closed-set VQA pairs along with test WSIs, aiming to evaluate various aspects of model performance. As shown in Fig. 3 (B), to improve the quality of the testing benchmarks, we employ four advanced large language models, including GPT-4 [2], InternLM2-Chat-7B [6], Qwen-7B-Chat [5], and DeepSeek-7B-Chat, to filter closed-set VQAs by predicting answers based solely on the question text. Any questions for which at least three of these models provided correct answers are subsequently excluded. Following this automated filtering, five expert pathologists are invited to review and amend the remaining questions. The review process is guided by the following criteria: (1) Whether the correct answer necessitates image interpretation; (2) Whether the question and its corresponding answer are logically and coherently structured; and (3) Whether the question aligns appropriately with the designated broad and narrow categories. VQA pairs failing to meet these criteria are excluded by the pathologists. Consequently, the SlideBench-VQA (TCGA) comprises 7,827 VQAs across 13 categories, with some examples illustrated in Fig. 3 C. Additionally, we incorporate the in-the-wild Early Breast Cancer Core-Needle Biopsy (BCNB) WSI dataset [30], which encompasses a diverse patient population and a variety of clinical task labels, to enhance the test set benchmark and more comprehensively assess the model’s generalization capabilities. In detail, we convert the BCNB dataset into a VQA format by rephrasing the classification objectives into a specific template as questions, while transforming the original multi-class labels into selectable options, and integrating it into SlideBench as an external subset, named SlideBench-VQA (BCNB). This dataset comprises 7,247 VQA pairs from 1,058 patients, specifically designed to evaluate SlideChat’s zero-shot generalization capability across 7 distinct classification tasks.